[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Sep 22, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Sep 22 20:36:12 UTC 2017


HSU Shift Log for Fri Sep 22 2017 16:36:01 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues: Craig helped us port AL98 data back into AL14, helping us  
restart Lee.

N-AWIPS issues: Had a strange hiccup on the Atlantic workstation (ridges)  
where NMAP windows were not loading after clicking the nmap2 button. After  
about 2 minutes, the windows then appeared.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Resubmitted meso sector request for Maria, running  
from 22/12Z to 23/06Z
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170922/d2352bd8/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list